At 1000 K, Kp = 19.9 for the following reaction:

Fe2O3(s)+3CO(g)⇌2Fe(s)+3CO2(g)

What are the equilibrium partial pressures of CO and CO2 if CO is the only gas present initially, at a partial pressure of 0.982 atm ?
Answer #1

Kp = [CO2]^3/[CO]^3

initial [CO] = 0.982 atm

let x be the change in pressure at equilibrium

19.9 = (3x^2)^3/(0.982 - 3x)^3

(0.982 - 3x) (0.982 - 3x) (0.982 - 3x)

0.964 - 5.892x + 9x^2

0.9466 - 2.892x - 5.786x + 17.676x^2 + 8.838x^2 - 27x^3 = 27x^3

54x^3 - 26.514x^2 + 8.678x - 0.9466 = 0

x = 0.163 atm

So the equilibrium concentration for,

[CO] = 0.982 - 3 x 0.163 = 0.493 atm

[CO2] =   0.489 atm